# 1.27in1.27in1in1in

### v1

@@ 2.4. Curve structures
[…]
@@
[…]
@@ 3. Interior special points
We put ourselves in the setup of the previous lemma. We call the point $p$ the interior special point of the curve $D$. A point in $D$ is special if it is either special interior or maps to one of the triple points under $\nu$.
@@ 3.1. Rulings on combinatorial K3 surfaces of class $T$
From the definition of $\ell$ we see that $D.L=2$, hence Riemann–Roch tells us that $\chi(L)=2$. Let $F_{0}$ be the effective Cartier divisor which is a linear combination of the $C_{i}$ representing $\ell$. Then $F_{0}\in|L|$, in particular, $L$ is effective. From Serre duality we infer that $H^{2}({\mathfrak{Y}}_{1},L)=0$ for otherwise $K_{{\mathfrak{Y}}_{1}}=K_{{\mathfrak{Y}}_{1}}-L+L$ were effective. Recall from [HL22, Construction 2.25] that ${\mathfrak{Y}}_{1}$ is obtained by contracting three curves in the anticanonical cycle of the $(1,5,1,3)$-blow-up [^5] An $n$-blow-up is a certain iterated blow-up depending on a tupel $n$ of nonnegative integers. For the precise definition, we refer to [HL22, Definition 2.24]. ${\widetilde{{\mathfrak{Y}}}}_{1}$ of ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ in certain points $(p_{1},p_{2},p_{3},p_{4})$ of its toric boundary. The exceptional divisor of the last blow-up in $p_{4}$ is a $(-1)$-curve on ${\widetilde{{\mathfrak{Y}}}}_{1}$ that becomes a curve $F$ of square zero on ${\mathfrak{Y}}_{1}$. By construction, its intersection numbers with the $C_{i}$ and with $D$ coincide with those of $L$, hence $F\in|L|$. In particular, the general member of $|L|$ is integral. Then, $L^{2}=0$ implies that $h^{0}(L)=2$, that $|L|$ is base point free, and hence that $L$ is nef. The remaining claims follow from the intersection number computation
$L.C_{i}=\begin{cases}1&\textrm{ for }i=1\\[5.0pt] 0&\textrm{ otherwise.}\end{cases}$
Note that the presence of another singular fiber would increase the Picard number. ∎
@@
Let $Y\in{\operatorname{CK3}_{2}}$ be a surface of class $T$ and let $Y_{T}\dashrightarrow Y$ be a composition of type I flops. Let $S\subset Y$ be a nonspecial component and consider the rational map $\phi:{\mathfrak{Y}}_{1}\dashrightarrow S$ given by the restriction of the above composition. If $S$ has Picard rank $\rho(S)\geq 2$, there is a unique fibration $S\xrightarrow{\ \ }{\mathbb{P}}^{1}$ that corresponds to the one from Lemma 3.3 under $\phi$. This fibration has a unique singular fiber.
@@
Using the behavior of curve structures under type I flops, see Section 5.1, we deduce that either $\phi$ or $\phi^{-1}$ is a morphism. In the latter case, the fibration is obtained by composition with this morphism, so let us assume that $\phi:{\mathfrak{Y}}_{1}\xrightarrow{\ \ }S$ is a morphism. By the enumeration of models [HL22, Section 7.2], [DSHL25] and the description of the fibration in Lemma 3.3, all curves contracted by $\phi$ are contained in fibers of ${\mathfrak{Y}}_{1}\xrightarrow{\ \ }{\mathbb{P}}^{1}$, so the fibration descends along such a fibration. The last claim is immediate if $\phi$ is regular. If $\phi^{-1}$ is regular, then one simply notes that $\phi^{-1}$ is a blow-up of ${\mathfrak{Y}}_{1}$ in the unique singular fiber of the fibration ${\mathfrak{Y}}_{1}\xrightarrow{\ \ }{\mathbb{P}}^{1}$. ∎
@@
The linear system $|L|$ is base point free by Corollary 3.5. Thus, point $p$ lies in a unique fiber of the corresponding morphism $S\xrightarrow{\ \ }{\mathbb{P}}^{1}$. It remains to show that this fiber is irreducible. If $S={\mathfrak{Y}}_{1}$, this is clear since there is only one singular fiber, this fiber entirely consists of curves in the curve structure, and none of the curves in the curve structure meets the node of the anticanonical curve. All other surfaces $S$ are either blow-ups or blow-downs of ${\mathfrak{Y}}_{1}$ where the center of the blow-up is always in the singular fiber. ∎
@@
By the enumeration of models, the only possibility for a surface $S$ as in Corollary 3.5 with Picard number $1$ is the following. With the curves in the curve structure numbered as in Figure 6, contract successively the curves $C_{8}$, $C_{7}$, $C_{6}$, $C_{5}$, $C_{4}$, $C_{3}$, $C_{2}$, $C_{1}$ (respectively their strict transforms). In this case, $S{\ \cong\ }{\mathbb{P}}^{2}$ and the anticanonical curve is a nodal cubic curve. The curve structure is given by one vertex, which is represented by a line. The final contraction, call it ${\widetilde{S}}\xrightarrow{\ \ }S$, is shown in Figure 7 at the level of curve structures. Lines passing through the interior special point on $S$ form a one-dimensional linear system, and there is a unique such line $F$ passing through the node of the anticanonical curve. The figure illustrates that this linear system is exactly the image of the linear system of the 00-curve on $S_{1}$ and the curve from Corollary 3.6 is exactly the strict transform ${\widetilde{F}}$ of $F$.
0011-1- 11111
@@
Let $S$ be a nonspecial component of a combinatorial K3 surface of class $T$ and let $p\in S$ be the node of the anticanonical curve $D$. Let $\pi:{\widetilde{S}}\xrightarrow{\ \ }S$ be the blow-up in the node and let $E:=\pi^{-1}(p)$ be the exceptional curve. Then we define an interior special point ${\widetilde{p}}$ of $E$ as follows. If $\rho(S)\geq 2$, there is a unique irreducible square zero curve $F$ in the linear system $|L|$ passing through $p$ by Corollary 3.6. We take ${\widetilde{p}}$ to be the unique intersection point $E\cap{\widetilde{F}}$ where ${\widetilde{F}}$ is the strict transform of $F$. Note that ${\widetilde{p}}$ is indeed interior for if ${\widetilde{F}}$ met the singular points of $\pi^{-1}(D)$, the curve $F$ would have to meet some branch of the node with multiplicity $\geq 2$, hence $F.D\geq 3$ contradicting the adjunction formula. If $\rho(S)=1$, we are in the situation of Example 3.7, in particular, $S{\ \cong\ }{\mathbb{P}}^{2}$. Here, we take ${\widetilde{p}}$ to be the unique intersection point $E\cap{\widetilde{F}}$ where ${\widetilde{F}}$ is the strict transform of the line $F$ joining the interior special point of $S$ with the node of the anticanonical curve.
Note that if ${\widetilde{S}}$ is an irreducible component of a combinatorial K3 surface (necessarily of class $P$), then $\pi^{*}D={\widetilde{D}}+E$ is the anticanonical cycle where ${\widetilde{D}}$ is the strict transform of the anticanonical curve $D$. Then the point ${\widetilde{p}}$ is indeed an interior special point according to our previous definition (Definition 3.2): it is the intersection of the $(-1)$-curve ${\widetilde{F}}$ with $E$ and ${\widetilde{F}}$ is in the curve structure by Corollary 5.5.
@@ 3.2. Rulings on combinatorial K3 surfaces of class $P$
Alternatively, we can successively blow down the curves $C_{1},C_{2},C_{3},C_{7},C_{6},C_{5}$. Then we are left with a smooth surface of Picard rank $2$. The images of $C_{4},C_{8}$ (which we denote by the same symbol) are curves of square 00 respectively $-2$ and $|C_{4}|$ gives a ruling with only smooth fibers. Since $C_{8}$ is still a section of square $-2$, we infer from the classification of surfaces that the contracted surface must be the second Hirzebruch surface ${\mathbb{F}}_{2}$. We obtain the sought-for fibration as the composite map $Y\rightarrow{\mathbb{F}}_{2}\rightarrow{\mathbb{P}}^{1}$. ∎
@@
$\ell^{\prime}=(0,0,1,2,2,2,2,1)\quad\textrm{ and }\quad\ell^{\prime\prime}=(2,2,2,2,1,0,0,1).$
@@
Let $S$ be a component of a combinatorial K3 surface of class $P$ and suppose that $C_{8}$ is not contracted under the rational map $\phi:{\mathfrak{Y}}_{2}\dashrightarrow S$. Then, for each triple point $p$ on $S$, there is a unique irreducible (square zero) curve in the linear system of $L$ passing through $p$.∎
@@
Note that if ${\widetilde{S}}$ is an irreducible component of a combinatorial K3 surface (necessarily the special component of a surface of class $T$), then $\pi^{*}(D_{1}+D_{2})={\widetilde{D}}_{1}+{\widetilde{D}}_{2}+E_{1}+E_{2}$ is the anticanonical cycle where ${\widetilde{D}}_{i}$ is the strict transform of $D_{i}$. Then the point ${\widetilde{p}}_{i}$ is indeed an interior special point according to Definition 3.2: it is the intersection of the $(-1)$-curve ${\widetilde{F}}_{i}$ with $E_{i}$ and ${\widetilde{F}}_{i}$ is in the curve structure by Corollary 5.5.
@@ 4. Type II flops
4.1. Type II flops of combinatorial K3 surfaces of degree $2$
Let us consider a combinatorial $K3$ surface and a smooth component $C=D_{ij}$ of the double locus of $Y$. Here, we adopt the notation of Section 2.2. The preimage of $C$ under the normalization map $\nu:Y^{\nu}\xrightarrow{\ \ }Y$ contains exactly two one-dimensional irreducible components and we assume that these are $(-1)$-curves in the respective normalizations of the components. The case $i=j$ is allowed, in which case $Y$ is a class $T$ surface and the preimage of $C=D_{ii}$ consists of two smooth disjoint $(-1)$-curves $D^{1}_{ii}\cup D^{2}_{ii}$. Otherwise, $Y$ is necessarily a class $P$ surface.
@@ 4.1. Type II flops of combinatorial K3 surfaces of degree $2$
We shall first treat the case $i\neq j$, which means that $Y$ is of class $P$. In this case, the curve $C$ intersects the third component $Y_{k}$ in two points $P^{1}_{k}$ and $P^{2}_{k}$. In this case we proceed as follows. We contract the $(-1)$ curves $D_{ij}\subset Y^{\nu}_{i}=Y_{i}$ and $D_{ji}\subset Y^{\nu}_{j}=Y_{j}$ to obtain smooth surfaces $Y^{\prime\nu}_{i}$ and $Y^{\prime\nu}_{j}$. The anticanonical cycles of these surfaces are now nodal rational curves. The component $Y_{k}=Y^{\nu}_{k}$ is also smooth and we blow it up in $P^{1}_{k}$ and $P^{2}_{k}$ to obtain a smooth surface $Y^{\prime\nu}_{k}$ with exceptional curves $E_{1}$ and $E_{2}$. The anticanonical cycle of this surface now consists of the cycle of four smooth components $D^{\prime\nu}_{ki}+E_{1}+D^{\prime\nu}_{kj}+E_{2}$ where $D^{\prime\nu}_{ki}$ and $D^{\prime\nu}_{kj}$ are the strict transforms of $D_{ki}$ and $D_{kj}$. The curves $D^{\prime\nu}_{ki}$ and $D^{\prime\nu}_{kj}$ have three special points, namely the singular points of the anticanonical cycle and the preimages of the interior special points. These are matched naturally to the three special points on $D^{\prime\nu}_{ik}$ and $D^{\prime\nu}_{jk}$. We now glue the two disjoint $(-1)$-curves $E_{1}$ and $E_{2}$ using the two singular points of the singular locus of the anticanonical cycle and the special interior points on $E_{1}$ and $E_{2}$ given by Definition 3.14. In this way we obtain a singular surface $Y^{\prime}_{k}$ and the images $D^{\prime}_{ki}$ and $D^{\prime}_{kj}$ of $D^{\prime\nu}_{ki}$ and $D^{\prime\nu}_{kj}$ become nodal curves. We now glue $Y^{\prime}_{i}$ and $Y^{\prime}_{k}$ along the nodal curves $D^{\prime}_{ik}$ and $D^{\prime}_{ki}$, thereby using the three special points on the normalizations of these curves (which become a node and an interior special point). Similarly, we glue $Y^{\prime}_{j}$ and $Y^{\prime}_{k}$. The result is a combinatorial K3 surface $Y^{\prime}$. Our matching of the special points ensures that the Carlson map is trivial and hence that $Y^{\prime}$ is maximal.
@@
We shall now consider the case $i=j$. Then we are in class $T$ with $Y_{i}$ being the special singular component. On the normalization $Y^{\nu}_{i}$ the preimage of $C=D_{ii}$ is a disjoint union $D^{1}_{ii}+D^{2}_{ii}$ of $(-1)$-curves. The anticanonical cycle is $D^{\nu}_{ij}+D^{1}_{ii}+D^{\nu}_{ik}+D^{2}_{ii}$. We blow down $D^{1}_{ii}$ and $D^{2}_{ii}$ and obtain a smooth surface $Y^{\prime}_{2}$ with anticanonal cycle $D^{\prime}_{ij}+D^{\prime}_{ik}$. Note that we have three special points on each of these curves. The curve $C$ intersects the two smooth components $Y_{j}$ and $Y_{k}$ in points $P_{j}$ and $P_{k}$ respectively. Blowing these up gives smooth surfaces $Y^{\prime}_{j}$ and $Y^{\prime}_{k}$ whose anticanocial cycles have two components $D^{\prime}_{ji}+E_{i}$ and $D^{\prime}_{ki}+E_{k}$ where $E_{j}$ and $E_{k}$ are the newly introduced exceptional curves. In this we we obtain the union $Y^{\prime\nu}=\cup_{\ell}Y^{\prime}_{\ell}$. We want to glue these to a singular surface $Y^{\prime}$. For this we identify $D^{\prime}_{ij}$ and $D^{\prime}_{ji}$ as well as $D^{\prime}_{ik}$ and $D^{\prime}_{ki}$ using the special points we have on these curves. Finally, we glue $E_{j}$ and $E_{k}$. For this we use the singular points of the anticanonical cycle and the new interior special points given by Definition 3.8. Again, the surface $Y^{\prime}$ has trivial Carlson homomorphism $c_{Y^{\prime}}=1$ and $Y^{\prime}$ is thus maximal.
@@
[…]
@@
One can generalize these definitions to arbitrary degree, but this requires a further analysis of maximal gluings. Abstractly, this is a consequence of the proof of [HL22, Proposition 2.23]. There it is shown that maximality is preserved under type I and II flops as the Carlson map pulls back to the Carlson map, see the arXiv version [AE21, Proposition 5.18]

### v2

@@ 2.4. Curve structures
A curve structure is a combinatorial object associated to a combinatorial K3 of degree $2$. We have met first instances in Example 2.15, here we define them for arbitrary components of surfaces in $\operatorname{CK3}_{2}$}. Recall the following result from [HL22, Corollary 6.32].
@@
of type I flops where $G\in\{P,T\}$. In particular, ${\operatorname{CK3}_{2}}(G)$ is the subclass of $\operatorname{CK3}_{2}$} of surfaces that can be obtained from $Y_{G}$ by a sequence of type I flops.
@@ 3. Interior special points
We put ourselves in the setup of the previous lemma. We call the point $p$ the interior special point of the curve $D$. A point in $D$ is special if it is either interior special or maps to one of the triple points under $\nu$.
@@ 3.1. Rulings on combinatorial K3 surfaces of class $T$
This can be read off from the description in Example 2.17. The sought-for fibration is given by pulling back the linear system of lines through the chosen inflection point of the nodal cubic in ${\mathbb{P}}^{2}$. Indeed, $C_{1}$ is the exceptional divisor of the first blow-up (and thus a section) and $\ell$ corresponds to the unique singular fiber which is the pullback of the flex line. ∎
@@
Let $Y\in{\operatorname{CK3}_{2}}$ be a surface of class $T$ and let $Y_{T}\dashrightarrow Y$ be a composition of type I flops. Let $S\subset Y$ be a nonspecial component and consider the rational map $\phi:{\mathfrak{Y}}_{1}\dashrightarrow S$ given by the restriction of the above composition. If $S$ has Picard rank $\rho(S)\geq 2$, there is a unique fibration $S\xrightarrow{\ \ }{\mathbb{P}}^{1}$ that corresponds to the one from Lemma 3.3 under $\phi$. This fibration has a unique singular fiber if $\rho(S)\geq 3$ and is smooth otherwise.
@@
From the description in Example 2.17, using the assumption $\rho(S)\geq 2$, we see that $S$ is obtained from ${\mathfrak{Y}}_{1}$ by blow-ups or blow-downs in the unique singular fiber of the fibration ${\mathfrak{Y}}_{1}\xrightarrow{\ \ }{\mathbb{P}}^{1}$. This proves the existence of the fibration, which by construction has at most one singular fiber. The claim follows. ∎
@@
The linear system $|L|$ is base point free by Corollary 3.5. Thus, the point $p$ lies in a unique fiber of the corresponding morphism $S\xrightarrow{\ \ }{\mathbb{P}}^{1}$. It remains to show that this fiber is irreducible. This follows from the description in Example 2.17, since the only (possibly) singular fiber is the preimage of the flex line while the fiber through $p$ is the strict transform of the line in ${\mathbb{P}}^{2}$ through the node and the chosen inflection point. ∎
@@
[…]
01-111
@@
[…]
Note that if $\widetilde{S}$} is an irreducible component of a combinatorial K3 surface (necessarily of class $P$), then $\pi^{*}D={\widetilde{D}}+E$ is the anticanonical cycle where $\widetilde{D}$} is the strict transform of the anticanonical curve $D$. Then the point $\widetilde{p}$} is indeed an interior special point according to our previous definition (Definition 3.2): it is the intersection of the (1)(-1)-curve $\widetilde{F}$} with $E$ and $\widetilde{F}$} is in the curve structure by Corollary 5.5.
@@ 3.2. Rulings on combinatorial K3 surfaces of class $P$
Alternatively, we can successively blow down the curves $C_{1},C_{2},C_{3},C_{7},C_{6},C_{5}$. Then we are left with a smooth surface of Picard rank $2$. The images of $C_{4},C_{8}$ (which we denote by the same symbol) are curves of square 0 respectively 2-2 and $|C_{4}|$ gives a ruling with only smooth fibers. Since $C_{8}$ is still a section of square 2-2, we infer from the classification of surfaces that the contracted surface must be the second Hirzebruch surface ${\mathbb{F}}_{2}$. We obtain the sought-for fibration as the composite map $Y\rightarrow{\mathbb{F}}_{2}\rightarrow{\mathbb{P}}^{1}$. ∎
@@
=(0,0,1,2,2,2,2,1) and $\ell^{\prime}=(0,0,1,2,2,2,2,1)\quad\textrm${ and $\quad\ell^{\prime\prime}$=(2,2,2,2,1,0,0,1).
@@
Let $S$ be a component of a combinatorial K3 surface of class $P$ and suppose that $C_{8}$ is not contracted under the rational map $\phi:{\mathfrak{Y}}_{2}\dashrightarrow S$. Then, for each triple point $p$ on $S$, there is a unique irreducible (square zero) curve in the linear system of $L$ passing through pp.∎
@@
Note that if $\widetilde{S}$} is an irreducible component of a combinatorial K3 surface (necessarily the special component of a surface of class $T$), then $\pi^{*}(D_{1}+D_{2})={\widetilde{D}}_{1}+{\widetilde{D}}_{2}+E_{1}+E_{2}$ is the anticanonical cycle where ${\widetilde{D}}_{i}$ is the strict transform of $D_{i}$. Then the point ${\widetilde{p}}_{i}$ is indeed an interior special point according to Definition 3.2: it is the intersection of the (1)(-1)-curve ${\widetilde{F}}_{i}$ with $E_{i}$ and ${\widetilde{F}}_{i}$ is in the curve structure by Corollary 5.5.
@@ 4. Type II flops
4.1. Type II flops of combinatorial K3 surfaces of degree 22
Let us consider a combinatorial $K3$ surface and a smooth component $C=D_{ij}$ of the double locus of $Y$. Here, we adopt the notation of Section 2.2. The preimage of $C$ under the normalization map $\nu:Y^{\nu}\xrightarrow{\ \ }Y$ contains exactly two one-dimensional irreducible components and we assume that these are (1)(-1)-curves in the respective normalizations of the components. The case $i=j$ is allowed, in which case $Y$ is a class $T$ surface and the preimage of $C=D_{ii}$ consists of two smooth disjoint (1)(-1)-curves $D^{1}_{ii}\cup D^{2}_{ii}$. Otherwise, $Y$ is necessarily a class $P$ surface.
@@ 4.1. Type II flops of combinatorial K3 surfaces of degree 22
We shall first treat the case $i\neq j$, which means that $Y$ is of class $P$. In this case, the curve $C$ intersects the third component $Y_{k}$ in two points $P^{1}_{k}$ and $P^{2}_{k}$. In this case we proceed as follows. We contract the (1)(-1) curves $D_{ij}\subset Y^{\nu}_{i}=Y_{i}$ and $D_{ji}\subset Y^{\nu}_{j}=Y_{j}$ to obtain smooth surfaces $Y^{\prime\nu}_{i}$ and $Y^{\prime\nu}_{j}$. The anticanonical cycles of these surfaces are now nodal rational curves. The component $Y_{k}=Y^{\nu}_{k}$ is also smooth and we blow it up in $P^{1}_{k}$ and $P^{2}_{k}$ to obtain a smooth surface $Y^{\prime\nu}_{k}$ with exceptional curves $E_{1}$ and $E_{2}$. The anticanonical cycle of this surface now consists of the cycle of four smooth components $D^{\prime\nu}_{ki}+E_{1}+D^{\prime\nu}_{kj}+E_{2}$ where $D^{\prime\nu}_{ki}$ and $D^{\prime\nu}_{kj}$ are the strict transforms of $D_{ki}$ and $D_{kj}$. The curves $D^{\prime\nu}_{ki}$ and $D^{\prime\nu}_{kj}$ have three special points, namely the singular points of the anticanonical cycle and the preimages of the interior special points. These are matched naturally to the three special points on $D^{\prime\nu}_{ik}$ and $D^{\prime\nu}_{jk}$. We now glue the two disjoint (1)(-1)-curves $E_{1}$ and $E_{2}$ using the two singular points of the singular locus of the anticanonical cycle and the special interior points on $E_{1}$ and $E_{2}$ given by Definition 3.14. In this way we obtain a singular surface $Y^{\prime}_{k}$ and the images $D^{\prime}_{ki}$ and $D^{\prime}_{kj}$ of $D^{\prime\nu}_{ki}$ and $D^{\prime\nu}_{kj}$ become nodal curves. We now glue $Y^{\prime}_{i}$ and $Y^{\prime}_{k}$ along the nodal curves $D^{\prime}_{ik}$ and $D^{\prime}_{ki}$, thereby using the three special points on the normalizations of these curves (which become a node and an interior special point). Similarly, we glue $Y^{\prime}_{j}$ and $Y^{\prime}_{k}$. The result is a combinatorial K3 surface $Y^{\prime}$. Our matching of the special points ensures that the Carlson map is trivial and hence that $Y^{\prime}$ is maximal.
@@
We shall now consider the case $i=j$. Then we are in class $T$ with $Y_{i}$ being the special singular component. On the normalization $Y^{\nu}_{i}$ the preimage of $C=D_{ii}$ is a disjoint union $D^{1}_{ii}+D^{2}_{ii}$ of (1)(-1)-curves. The anticanonical cycle is $D^{\nu}_{ij}+D^{1}_{ii}+D^{\nu}_{ik}+D^{2}_{ii}$. We blow down $D^{1}_{ii}$ and $D^{2}_{ii}$ and obtain a smooth surface $Y^{\prime}_{2}$ with anticanonal cycle $D^{\prime}_{ij}+D^{\prime}_{ik}$. Note that we have three special points on each of these curves. The curve $C$ intersects the two smooth components $Y_{j}$ and $Y_{k}$ in points $P_{j}$ and $P_{k}$ respectively. Blowing these up gives smooth surfaces $Y^{\prime}_{j}$ and $Y^{\prime}_{k}$ whose anticanocial cycles have two components $D^{\prime}_{ji}+E_{i}$ and $D^{\prime}_{ki}+E_{k}$ where $E_{j}$ and $E_{k}$ are the newly introduced exceptional curves. In this we we obtain the union $Y^{\prime\nu}=\cup_{\ell}Y^{\prime}_{\ell}$. We want to glue these to a singular surface $Y^{\prime}$. For this we identify $D^{\prime}_{ij}$ and $D^{\prime}_{ji}$ as well as $D^{\prime}_{ik}$ and $D^{\prime}_{ki}$ using the special points we have on these curves. Finally, we glue $E_{j}$ and $E_{k}$. For this we use the singular points of the anticanonical cycle and the new interior special points given by Definition 3.8. Again, the surface $Y^{\prime}$ has trivial Carlson homomorphism $c_{Y^{\prime}}=1$ and $Y^{\prime}$ is thus maximal.
@@
$D_{ji}$$D_{ij}$$P^{\nu}_{k}$$P^{\nu}_{\ell}$$Y_{j}^{\nu}$$Y_{i}^{\nu}$$Y^{\nu}_{k}$$Y^{\nu}_{\ell}$$D^{\prime}_{k\ell}$$D^{\prime}_{\ell k}$$P^{\prime\nu}_{i}$$P^{\prime\nu}_{j}$$Y_{k}^{\prime\nu}$$Y_{\ell}^{\prime\nu}$$Y^{\prime\nu}_{j}$$Y^{\prime\nu}_{i}$$\phi^{\nu}$
@@
One can generalize these definitions to arbitrary degree, but this requires a further analysis of maximal gluings. Abstractly, this is a consequence of the proof of [HL22, Proposition 2.23]. There it is shown that maximality is preserved under type I and II flops as the Carlson map pulls back to the Carlson map; see the arXiv preprint [AE21, Proposition 5.18].